princeevee:

A coin is flipped four times. X is the number of times it will land with a head up. Which statement is true about this situation? The variable X has a binomial distribution. P(success)=0.5P(success)=0.5; number trials = 4 The variable X does not have a binomial distribution because P(success) is not constant. The variable X does not have a binomial distribution because there are more than two possible outcomes. The variable X has a binomial distribution. P(success)=0.25; number of trials = 4

Vocaloid:

hint: what is the probability of heads on ONE coin?

princeevee:

0.5?

Vocaloid:

good so P(success) = 0.5 since there are two outcomes, this is a binomial distribution with number of trials = 4 so option A must be your answer

princeevee:

can you help with more please?

Vocaloid:

ok

princeevee:

For a certain type of hay fever, Medicine H has a 30% probability of working. In which distributions does the variable X have a binomial distribution? Select each correct answer. When the medicine is tried with two patients, X is the number of doses each patient needs to take. When the medicine is tried with six patients, X is the number of patients for whom the medicine does not work. When the medicine is tried with two patients, X is the number of patients for whom the medicine worked. When the medicine is tried with six patients, X is the number of patients for whom the medicine worked.

Vocaloid:

I believe it would be everything except the first option since X = the number of patients (success or fail)

Vocaloid:

if that doesn't work then I would only pick the last two

princeevee:

1 attachment
Vocaloid:

hint: if P(success) * P(success) = 0.36 then what is P(success)?

princeevee:

9?

princeevee:

no wait

Vocaloid:

what is the square root of 0.36?

princeevee:

0.6

Vocaloid:

good, that's your answer

princeevee:

1 attachment
Vocaloid:

hint: what is the probability of landing on A or B?

princeevee:

1/3

Vocaloid:

in probability "or" implies addition probability of A + probability of B = ?

princeevee:

2/3?

Vocaloid:

good, and if we square that (for X = 2) we get?

princeevee:

B?

Vocaloid:

what is (2/3)*(2/3) ?

princeevee:

1

Vocaloid:

what is 2*2?

princeevee:

4

Vocaloid:

good, what is 3*3?

princeevee:

9

Vocaloid:

good, so (2/3)*(2/3) = 4/9 so if X = 2 gives us a probability of 4/9 that means only option #1 (top left) works

princeevee:

The probability that traffic lights on a certain road will be green is 0.6. When a driver on that road approaches two traffic lights in a row, X is the number of traffic lights that are green. What is P(X=1)? Enter your answer, as a decimal,

Vocaloid:

if 0.6 is the probability of green, what is the probability of NOT green?

princeevee:

0.4?

Vocaloid:

good so we have two possibilities green * not green not green* green so that gives us (0.6)(0.4) + (0.4)(0.6) = ?

princeevee:

0.48

Vocaloid:

good that should be your answer

princeevee:

thanks!